OpenStudy (anonymous):

plz help me......... (x^2-2xy-y^2)dx=(x+y)^2dy

OpenStudy (anonymous):

can any one get the ans?????????

OpenStudy (anonymous):

yup i m trying to slove

OpenStudy (anonymous):

\[y=vx\]Use this substitution.

OpenStudy (anonymous):

\[\frac{dy}{dx} = v + x\frac{dv}{dx}\]

OpenStudy (anonymous):

kk let me try ya i got u.........

OpenStudy (anonymous):

is this a question to prove?

OpenStudy (anonymous):

\[(x^2-2xy-y^2)dx=(x+y)^2dy \\ \frac {dy}{dx}=\frac {x^2-2xy-y^2} {(x+y)^2}=\frac{(x-y)^2}{(x+y)^2} \]

OpenStudy (anonymous):

bt (x-y)^2 == x^2-2xy +y^2 na?????

OpenStudy (anonymous):

nope this que for solve

OpenStudy (anonymous):

what are we trying to find?